OpenStudy (anonymous):

ANSWER FAST PLEASE 29.) 4k-1≥-2 30.) 6(c-1)<-18 31.)3t>5t+12 32) -6/7y-6≥42 33. 4+x/2>2x 34.) 3x+5≤2x-8 35.) 42w>(w+7)

OpenStudy (hoblos):

\[29) k \ge -1/4\] \[30) c<2\] \[31) t<-6\] \[32) y \le 56\] \[33) x<8/3\] \[34) x \le -13\]

OpenStudy (hoblos):

\[35) w> 7/41\]

OpenStudy (anonymous):

thank you thank you thank you!

OpenStudy (hoblos):

wlc wlc wlc lol